# A. CamillettiAI weather from Euro-Atlantic regimes

### v1

@@ Abstract
We present a non-linear AI-model designed to reconstruct monthly mean anomalies of the European temperature and precipitation based on the Euro-Atlantic Weather regimes (WR) indices. WR represent recurrent, quasi-stationary, and persistent states of the atmospheric circulation that exert considerable influence over the European weather, therefore offering an opportunity for sub-seasonal to seasonal forecasting. While much research has focused on studying the correlation and impacts of the WR on European weather, the estimation of ground-level climate variables, such as temperature and precipitation, from Euro-Atlantic WR remains largely unexplored and is currently limited to linear methods. The presented AI model can capture and introduce complex non-linearities in the relation between the WR indices, describing the state of the Euro-Atlantic atmospheric circulation and the corresponding surface temperature and precipitation anomalies in Europe. We discuss the AI-model performance in reconstructing the monthly mean two-meter temperature and total precipitation anomalies in the European winter and summer, also varying the number of WR used to describe the monthly atmospheric circulation. We assess the impact of errors on the WR indices in the reconstruction and show that a mean absolute relative error below 80% yields improved seasonal reconstruction compared to the ECMWF operational seasonal forecast system, SEAS5. As a demonstration of practical applicability, we evaluate the model using WR indices predicted by SEAS5, finding slightly better or comparable skill relative to the SEAS5 forecast itself. Our findings demonstrate that WR-based anomaly reconstruction, powered by AI tools, offers a promising pathway for sub-seasonal and seasonal forecasting.
keywords:
Euro-Atlantic regimes, seasonal prediction, artificial intelligence, circulation patterns, reconstruction
@@ 1 Introduction
Predicting atmospheric conditions for the upcoming season is crucial in various social and economic applications. Accurate seasonal forecasts can guide agricultural planning, allowing farmers to optimize crop selection, planting, and harvesting schedules, supporting food security and market stability [Ceglar and Toreti, 2021]. In the energy sector, such forecasts help anticipate demand fluctuations, aiding in efficient resource allocation and grid management [Dorrington et al., 2020, Buonocore et al., 2022, El-Azab et al., 2025]. Additionally, governments and disaster preparedness agencies can use reliable forecasts to mitigate the impacts of extreme weather events, such as droughts or floods, thereby reducing economic losses and protecting vulnerable communities [Alfieri et al., 2013, Samaniego et al., 2019]. Despite recent advances, forecasting European weather on long-term timescales remains challenging for both dynamical [Palmer et al., 2004, Mishra et al., 2019] and statistical methods [Krikken et al., 2024]. While the weather forecast of physics-based numerical weather prediction (NWP) is limited to about two weeks due to the intrinsic chaotic nature of the atmosphere [Lorenz, 1963], statistical methods can explicitly leverage slowly varying, persistent, and more predictable components of the Earth system. By Rossby wave propagation, these large-scale phenomena give rise to teleconnection patterns between distant regions of the Earth [Wallace and Gutzler, 1981, Ambrizzi et al., 1995], opening a window of opportunity to enlarge the forecast horizons. In the northern hemisphere, the main mode of atmospheric variability, represented by the North Atlantic Oscillation (NAO) [Hurrell et al., 2003], or the strongly related Artic Oscillation (AO), exerts influence over North America, Europe and part of Asia [Deser et al., 2017, Trigo et al., 2002], and is characterized by a typical persistence of 10 to 30 days [Feldstein, 2003, Rennert and Wallace, 2009, Roberts et al., 2023].
@@
The existence of persistent modes of the atmospheric state suggests that large-scale atmospheric circulation can be systematically decomposed into a finite number of dominant states [Rossby, 1940]. Weather regimes (WR), first introduced by Rex [1951], are recurrent, quasi-stationary, and persistent states of atmospheric circulation [Hochman et al., 2021]. In the Euro-Atlantic sector, the WR are strongly related to the NAO [Michelangeli et al., 1995] and are known to exert considerable influence over the European weather [van der Wiel et al., 2019, Trigo and DaCamara, 2000, Plaut and Simonnet, 2001, Hall and Hanna, 2018, Simpson et al., 2024] [^1] Some of the works mentioned here refer to different large-scale circulation patterns than WR, but share similar ideas and comparable results.. Their persistent nature [see e.g. Strommen et al., 2023, Mukhin et al., 2024b, for a related discussion], often influenced by teleconnection patterns [Roberts et al., 2023, Michel and Rivière, 2011], opens up the possibility of extending the forecast horizon in Europe beyond the limit posed by the chaotic nature of the atmosphere [Ferranti et al., 2015, 2018]. This justifies the increasing interest in assessing the capability of NWP [Ferranti et al., 2018, Büeler et al., 2021] and statistical models [Nielsen et al., 2022, Hall et al., 2019, Chattopadhyay et al., 2020, Bommer et al., 2025] to forecast the transitions between large-scale circulation regimes in the sub-seasonal and seasonal timescales. However, while much research has focused on the study of the correlation and the impacts of the WR on the European weather, the estimation of ground-level climate variables, such as temperature and precipitation, from Euro-Atlantic WR remains largely unexplored and limited to linear methods, e.g. considering the mean surface weather variables associated with the dominant regime [Madonna et al., 2021, Roberts et al., 2023]. Additionally, Gerighausen et al. [2024] highlight the importance of intra-regime variability and suggest the use of continuous WR indices to capture the regimes ‘subflavours’ in place of the categorical classification commonly found in the literature [Madonna et al., 2021, Grams et al., 2017, Plaut and Simonnet, 2001, Trigo and DaCamara, 2000].
@@
In this study, we present an Artificial Intelligence (AI) model designed to reconstruct anomalies in European temperature and precipitation based on knowledge of the WR indices. The model can capture and introduce complex non-linearities in the relationship between the WR indices, which describe the state of the Euro-Atlantic atmospheric circulation, and the corresponding ground temperature and precipitation anomalies in Europe. After presenting the datasets (Sec. 2) and defining the WR and the target fields (Sec. 3), we present the AI model in Sec. 4. We evaluate its accuracy in reconstructing seasonal mean temperature and total precipitation anomalies in winter and summer (Sec. 5.1), as well as the impact of changing the number of indices used to represent the atmospheric circulation (Sec. 5.2). Given the potential future applications in a real forecasting scenario, we discuss the effects of introducing errors in the WR indices, establishing a lower bound on the WR prediction accuracy required to outperform the European Centre for Medium-Range Weather Forecasts (ECMWF) operational seasonal forecast system, SEAS5 (Sec. 5.3). In other words, we utilize the presented AI model as a mapping from the WR indices to the anomalies of interest to address the following questions:
@@
We also test the accuracy of the AI model in reconstructing the targeted anomalies using the SEAS5 forecast of the winters and summers WR indices (Sec. 5.4) as a first evaluation of the model in a concrete forecasting scenario.
In summary, the AI model presented here offers a promising approach to AI-driven long-range weather forecasting by leveraging the predictability of large-scale atmospheric circulation and its relation with surface-level variables. The model demonstrates good reconstruction capability, as well as notable adaptability and stability in the presence of errors in the WR indices, underlining its potential for seasonal forecasting applications. While it is applied in this study to reconstruct two-meter temperature and total precipitation anomalies on a monthly timescale, the architecture is sufficiently flexible to be adapted to other regions and variables without modification. Additionally, the model’s relatively lightweight design enables efficient training, further supporting its practicality for broader implementation.
@@ 2.1 ERA5 reanalysis
ERA5, the fifth generation of ECMWF global reanalysis [Hersbach et al., 2020], is the current state-of-the-art reanalysis and the most widely used data set in climate sciences. The ERA5 reanalysis spans years from 1940 onward, with a latency of 5 days and a Cartesian grid resolution of 0.25°, corresponding to approximately 31 km horizontal resolution at the equator. Using a bilinear interpolation, we upscaled the ERA5 data from the original 0.25°×0.25°Cartesian grid to a 1°×1°Cartesian grid.
The geopotential height (GPH) is one of the most widely used variables for studying teleconnections and decomposing atmospheric circulation into a set of dominant modes. We use the daily mean GPH at 500 hPa (Z500) in the Euro-Atlantic region (80°W to 40°E, 30°N to 90°N) to compute the WR and the related WR indices (see Sec. 3.1) used as input in our model. The work aims to reconstruct the monthly mean two-meter temperature and total precipitation anomalies in Europe (i.e., within the domain 20°W to 40°E, 30°N to 70°N). We use the daily mean two-meter temperature and daily total precipitation to compute the respective daily anomalies, which are then aggregated monthly (see Sec. 3.3).
@@ 2.2 Numerical seasonal forecast: SEAS5
SEAS5 is the fifth-generation seasonal forecasting system developed by ECMWF [Johnson et al., 2019]. It provides global, long-range forecasts up to seven months ahead, using an advanced coupled ocean-atmosphere model. The SEAS5 forecast and hindcast of monthly aggregated GPH, two-meter temperature, and precipitation with 25 ensemble members were retrieved from the Copernicus Data Store (CDS) in the months from January 1981 to December 2024 on a 1°×1°regular grid over the Euro-Atlantic region previously defined.
@@ 3 Pre-processing
The following section details the pre-processing steps employed to derive the input and target variables for the AI-model. We then describe the model architecture, training and evaluation in Sec. 4.
3.1 Model’s input: weather regimes indices
Figure 1: Cluster mean of the $Z500$ field (black isocontours) and corresponding average $Z500$ anomalies in winter (December, January, February), associated with the seven year-round Euro-Atlantic WR. The average is performed in the 1981-2010 climatology. The ”No Regime” refers to the mean atmospheric circulation when there is no WR attribution according to the index life-cycle criteria [Grams et al., 2017].
There are different methods of extracting low-frequency modes of variability of the atmospheric circulation [refer to Hannachi et al., 2017, Schlef et al., 2019, Mukhin et al., 2024a, Spuler et al., 2024, for some examples]. In this work, we use the method presented in Grams et al. [2017], a natural extension of the procedure discussed in Ferranti et al. [2015]. The following briefly recap the procedure for obtaining the seven-year-round WR [Grams et al., 2017, Büeler et al., 2021].
@@ 3.1 Model’s input: weather regimes indices
    The calendar day climatology is obtained by applying a 15-day moving window centered on each calendar day (±$\pm~{}7$ days) in the climatology period;
@@
    The resulting field is 5 days low-pass filtered using a Savitzky–Golay filter;
@@
    A calendar day normalization is computed by applying to the $Z500$ calendar standard deviation a 15-day moving window centered on each calendar day (±$\pm~{}7$ days) in the climatology period, and then taking the spatial average on the full domain;
@@
$I_{\rm wr}(t)=\frac{P_{\rm wr}(t)-\left\langle P_{\rm wr}(t)\right\rangle}{\sqrt{\left\langle\left(P_{\rm wr}(t)-\left\langle P_{\rm wr}(t)\right\rangle\right)^{2}\right\rangle}}~{},$ (1)
[…]
@@
The method outlined here for deriving the seven WR differs slightly from that used in Grams et al. [2017] and Büeler et al. [2021]. Key differences are detailed in Appendix B.
We are interested in quantifying the performance of the AI model in reconstructing the targeted anomalies by changing the number of WR used as input. Therefore, in addition to the seven weather regimes, we use the same procedure described before for the winter (December, January, February - DJF) and summer (June, July, August - JJA) seasons separately, changing the number of clusters to four. The four weather regimes obtained are the ZO, the GL, the EuBL, and the AR in winter (Fig. A.1), and the negative phase of the NAO (NAO-), a blocking state similar to the ScBL, the ZO (often referred as Atlantic Low [Cassou et al., 2005]) and the AT, in summer (Fig. A.2). We also compute the NAO index [Hurrell et al., 2003] separately for the summer and winter seasons, as the principal component (PC) of the first EOF. The analysis is performed on the same Euro-Atlantic region and the same climatology period used to compute the WR.
Finally, we aggregate the $I_{\rm wr}$ monthly. Indeed, we use the monthly average of the seven or four $I_{\rm wr}$ as an input variable in our AI model (see Sec. 4).
@@ 3.2 Perturbed indices
For every value of the MARE from 0 to 160%, we compute 50 random realization of $I_{\rm wr}^{\prime}$ sampling from $I_{\rm wr}\times N\left(0,\sqrt{\pi/2}~{}{\rm{MARE}}\right)$. Each realization is then passed as an input to the AI model (see Sec. 4) to obtain an estimate of the monthly average two-meter temperature and total precipitation in Europe for the months from January 2011 to December 2024 .

### v2

# AI reconstruction of European weather from the Euro-Atlantic regimes

@@ Abstract
We present a non-linear AI model designed to reconstruct monthly mean anomalies of the European temperature and precipitation based on the Euro-Atlantic Weather regimes (WR) indices. WR represent recurrent, quasi-stationary, and persistent states of the atmospheric circulation that exert considerable influence over the European weather, therefore offering an opportunity for sub-seasonal to seasonal forecasting. While much research has focused on studying the correlation and impacts of the WR on European weather, the estimation of ground-level meteorological variables, such as temperature and precipitation, from Euro-Atlantic WR remains largely unexplored and is currently limited to linear methods. The presented AI model can capture and introduce complex non-linearities in the relation between the WR indices, describing the state of the Euro-Atlantic atmospheric circulation and the corresponding surface temperature and precipitation anomalies in Europe. We discuss the AI-model performance in reconstructing the monthly mean two-meter temperature and total precipitation anomalies in the European winter and summer, also varying the number of WR used to describe the monthly atmospheric circulation. We assess the impact of errors on the WR indices in the reconstruction and show that a mean absolute relative error below 80% yields improved seasonal reconstruction compared to the ECMWF operational seasonal forecast system, SEAS5. As a demonstration of practical applicability, we evaluate the model using WR indices predicted by SEAS5, finding slightly better or comparable skill relative to the SEAS5 forecast itself. Our findings demonstrate that WR-based anomaly reconstruction, powered by AI tools, offers a promising pathway for sub-seasonal and seasonal forecasting.
Keywords Euro-Atlantic regimes, seasonal prediction, artificial intelligence, circulation patterns, reconstruction
@@ 1 Introduction
Predicting atmospheric conditions for the upcoming season is crucial in various social and economic applications. Accurate seasonal forecasts can guide agricultural planning, allowing farmers to optimize crop selection, planting, and harvesting schedules, supporting food security and market stability [ceglar_seasonal_2021]. In the energy sector, such forecasts help anticipate demand fluctuations, aiding in efficient resource allocation and grid management [Dorrington_2020, buonocore_inefficient_2022, el-azab_seasonal_2025]. Additionally, governments and disaster preparedness agencies can use reliable forecasts to mitigate the impacts of extreme weather events, such as droughts or floods, thereby reducing economic losses and protecting vulnerable communities [alfieri_glofas_2013, Samaniego_2019]. Despite recent advances, forecasting European weather on long-term timescales remains challenging for both dynamical [palmer_development_2004, mishra_multi-model_2019] and statistical methods [Krikken_2024]. While the weather forecast of physics-based numerical weather prediction (NWP) is limited to about two weeks due to the intrinsic chaotic nature of the atmosphere [lorenz_deterministic_1963], statistical methods can explicitly leverage slowly varying, persistent, and more predictable components of the Earth system. By Rossby wave propagation, these large-scale phenomena give rise to teleconnection patterns between distant regions of the Earth [wallace_teleconnections_1981, ambrizzi_rossby_1995], opening a window of opportunity to enlarge the forecast horizons. In the northern hemisphere, the main mode of atmospheric variability, represented by the North Atlantic Oscillation (NAO) [hurrell_overview_2003], or the strongly related Artic Oscillation (AO), exerts influence over North America, Europe and part of Asia [deser_role_2017, trigo_2002], and is characterized by a typical persistence of 10 to 30 days [feldstein_dynamics_2003, rennert_cross-frequency_2009, roberts_euro-atlantic_2023].
@@
The existence of persistent modes of the atmospheric state suggests that large-scale atmospheric circulation can be systematically decomposed into a finite number of dominant states [Rossby_1940]. Weather regimes (WR), first introduced by rex_effect_1951, are recurrent, quasi-stationary, and persistent states of atmospheric circulation [hochman_atlantic-european_2021]. WR may be defined separately for the summer and winter seasons, in which case four WR are optimal in the Euro-Atlantic sector [michelangeli_weather_1995], or over the entire year, in which case seven WR are optimal [grams_balancing_2017]. In the Euro-Atlantic sector, the WR are strongly related to the NAO [michelangeli_weather_1995] and are known to exert considerable influence over the European weather [van_der_wiel_influence_2019, trigo_circulation_2000, plaut_large-scale_2001, hall_north_2018, simpson_north_2024] [^1] Some of the works mentioned here refer to different large-scale circulation patterns than WR, but share similar ideas and comparable results.. Their persistent nature [see e.g. strommen_topological_2023, mukhin_metastability_2024, for a related discussion], often influenced by teleconnection patterns [roberts_euro-atlantic_2023, michel_link_2011], opens up the possibility of extending the forecast horizon in Europe beyond the limit posed by the chaotic nature of the atmosphere [ferranti_flow-dependent_2015, ferranti_how_2018]. This justifies the increasing interest in assessing the capability of NWP [ferranti_how_2018, bueler_year-round_2021] and Artificial Intelligence (AI) models based on Neural Networks (NN) architectures to forecast the transitions between large-scale circulation regimes in the sub-seasonal and seasonal timescales. In the sub-seasonal scale, deep learning models such as convolutional neural networks (CNNs) and recurrent neural networks (RNNs) have been successfully applied to enhance large-scale circulation regimes by capturing complex spatial and temporal dependencies in reanalysis data [chattopadhyay_predicting_2020, nielsen_forecasting_2022]. bommer_deep_2025 introduce advanced machine learning approaches to improve sub-seasonal predictions of the WR by leveraging teleconnection patterns, further enhancing forecast reliability and skill. On the seasonal scale, hall_complex_2019 use complex modeling techniques to address the data-scarce conditions typical of this scale, aiming to forecast the first three principal components derived from empirical orthogonal function (EOF) analysis in the North Atlantic region.
However, while much research has focused on the evaluation of the forecast and on the impacts of the WR on the European weather, the estimation of ground-level climate variables, such as temperature and precipitation, from Euro-Atlantic WR remains largely unexplored and limited to linear methods, e.g. considering the mean surface weather variables associated with the dominant regime [madonna_reconstructing_2021, roberts_euro-atlantic_2023]. Additionally, gerighausen_quantifying_2024 highlight the importance of intra-regime variability and suggest the use of continuous WR indices to capture the regimes ‘subflavours’ in place of the categorical classification commonly found in the literature [madonna_reconstructing_2021, grams_balancing_2017, plaut_large-scale_2001, trigo_circulation_2000].
@@
In this study, we present two AI models designed to reconstruct monthly anomalies in European temperature or precipitation, based on knowledge of the WR indices. The models can capture and introduce complex non-linearities in the relationship between the WR indices, which describe the state of the Euro-Atlantic atmospheric circulation, and the corresponding two-meter temperature and precipitation anomalies in Europe. After presenting the datasets (Sec. 2) and defining the WR and the target fields (Sec. 3), we present the AI model’s architecture in Sec. 4. We evaluate its accuracy in reconstructing monthly mean temperature and total precipitation anomalies in winter and summer (Sec. 5.1) and compare it with a standard baseline, namely the linear regression on individual WR (Sec. 5.2). We then discuss the impact of changing the number of indices used to represent the atmospheric circulation (Sec. 5.3). Given the potential future applications in a real forecasting scenario, we discuss the effects of introducing errors in the WR indices, establishing a lower bound on the WR prediction accuracy required to outperform the European Center for Medium-Range Weather Forecasts (ECMWF) operational seasonal forecast system, SEAS5 (Sec. 5.4). In other words, we utilize the presented AI models as a mapping from the WR indices to the anomalies of interest to address the following questions:
@@
We also test the accuracy of the AI models in reconstructing the targeted anomalies using the SEAS5 forecast of the winters and summers WR indices (Sec. 5.5) as a first evaluation of the models in a concrete forecasting scenario.
In summary, the AI models presented here offer a promising approach to AI-driven long-range weather forecasting by leveraging the predictability of large-scale atmospheric circulation and its relation with surface-level variables. The models demonstrate good reconstruction capability, as well as notable adaptability and stability in the presence of errors in the WR indices, underlining their potential for seasonal forecasting applications. While they are applied in this study to reconstruct two-meter temperature and total precipitation anomalies on a monthly timescale, the architecture is sufficiently flexible to be adapted to other regions and variables without modification. Additionally, the model’s relatively lightweight design enables efficient training (requiring a single GPU), further supporting its practicality for broader implementation.
@@ 2.1 ERA5 reanalysis
ERA5, the fifth generation of ECMWF global reanalysis [Hersbach2020], is the current state-of-the-art reanalysis and the most widely used data set in climate sciences. The ERA5 reanalysis spans years from 1940 onward, with a latency of 5 days and a Cartesian grid resolution of 0.25°, corresponding to approximately 31 km horizontal resolution at the equator. Using a bilinear interpolation, we upscaled the ERA5 data from the original 0.25°×0.25°Cartesian grid to a 1°×1°Cartesian grid.
The geopotential height (GPH) is one of the most widely used variables for studying teleconnections and decomposing atmospheric circulation into a set of dominant modes. We use the daily mean GPH at 500 hPa (Z500) in the Euro-Atlantic region (80°W to 40°E, 30°N to 90°N) to compute the WR and the related WR indices (see Sec. 3.1) used as input in our models. The work aims to reconstruct the monthly mean two-meter temperature and total precipitation anomalies in Europe (i.e., within the domain 20°W to 30°E, 35°N to 70°N). We use the daily mean two-meter temperature and daily total precipitation to compute the respective daily anomalies, which are then aggregated monthly (see Sec. 3.3).
@@ 2.2 Numerical seasonal forecast: SEAS5
SEAS5 is the fifth-generation seasonal forecasting system developed by ECMWF [johnson_seas5_2019]. It provides global, long-range forecasts up to seven months ahead, using an advanced coupled ocean-atmosphere model. The SEAS5 forecast and hindcast of monthly aggregated GPH at 500 hPa, two-meter temperature, and precipitation with 25 ensemble members were retrieved from the Copernicus Data Store (CDS) in the months from January 1981 to December 2024 on a 1°×1°regular grid over the Euro-Atlantic region previously defined.
@@ 3 Pre-processing
The following section details the pre-processing steps employed to derive the input and target variables for the AI models. We then describe the models’ architecture, training and evaluation in Sec. 4.
3.1 Input of the AI-model: weather regimes indices
Figure 1: Cluster mean of the $Z500$ field (black isocontours) and corresponding average $Z500$ anomalies in winter (December, January, February), associated with the seven year-round Euro-Atlantic WR. The average is performed in the 1981-2010 climatology. The “No Regime” refers to the mean atmospheric circulation when there is no WR attribution according to the index life-cycle criteria [grams_balancing_2017].
There are different methods of extracting low-frequency modes of variability of the atmospheric circulation [refer to hannachi_low-frequency_2017, schlef_atmospheric_2019, mukhin_revealing_2024, spuler_identifying_2024, for some examples]. In this work, we use the method presented in grams_balancing_2017, a natural extension of the procedure discussed in ferranti_flow-dependent_2015. The following briefly recaps the procedure for obtaining the seven year-round WR [grams_balancing_2017, bueler_year-round_2021].
@@ 3.1 Input of the AI-model: weather regimes indices
    The calendar day climatology is obtained by applying a 15-day moving window centered on each calendar day (± $\pm\penalty 10000\ 7$ days) in the climatology period;
@@
    The resulting field is 5 days low-pass filtered using a Savitzky–Golay filter [^2] In agreement with robertson2020 and lee2023, we did not find significant differences between the low-pass filtered and non-filtered WR.;
@@
    A calendar day normalization is computed by applying to the $Z500$ calendar standard deviation a 15-day moving window centered on each calendar day (± $\pm\penalty 10000\ 7$ days) in the climatology period, and then taking the spatial average on the full domain;
@@
$I_{\rm wr}(t)=\frac{P_{\rm wr}(t)-\left\langle P_{\rm wr}(t)\right\rangle}{\sqrt{\left\langle\left(P_{\rm wr}(t)-\left\langle P_{\rm wr}(t)\right\rangle\right)^{2}\right\rangle}}\penalty 10000\ ,$ (1)
where the averages $\langle\cdot\rangle$ are performed in the climatology period, and $P_{\rm wr}$ is the aforementioned projection [see michelangeli_weather_1995]:
@@
The method outlined here for deriving the seven WR differs slightly from that used in grams_balancing_2017 and bueler_year-round_2021. Key differences are detailed in Appendix B.
We are interested in quantifying the performance of the AI models in reconstructing the targeted anomalies by changing the number of WR used as input. Therefore, in addition to the seven weather regimes, we use the same procedure described before for the winter (December, January, February - DJF) and summer (June, July, August - JJA) seasons separately, changing the number of clusters to four. The four weather regimes obtained are the AT, the ScBL, the ZO, and the GL in winter (Fig. A.1), and in summer (Fig. A.2). We also compute the NAO index [hurrell_overview_2003] separately for the summer and winter seasons, as the principal component (PC) of the first EOF. The analysis is performed on the same Euro-Atlantic region and the same climatology period used to compute the WR.
Finally, we aggregate the $I_{\rm wr}$ on a monthly basis. Indeed, we use the monthly average of the seven or four $I_{\rm wr}$ as an input variable in our AI models (see Sec. 4).
@@ 3.2 Perturbed indices
For every value of the MARE from 0 to 160%, we compute 50 random realization of $I_{\rm wr}^{\prime}$ sampling from $I_{\rm wr}\times N\left(0,\sqrt{\pi/2}\penalty 10000\ {\rm{MARE}}\right)$. Each realization is then passed as an input to the AI models (see Sec. 4) to obtain an estimate of the monthly average two-meter temperature and total precipitation in Europe for the months from January 2011 to December 2024 .